OpenStudy (anonymous):

Any tips on how to do mental math? I was wondering if you geniuses have any technique of doing equation in mind...for example if someone asks me 199+5431 I need a sheet of paper and a pen to write it down...i just can't do it in my mind,even harder with multiplication...whats wrong? :(

OpenStudy (konradzuse):

1 9 9 5 4 3 1 9 + 1 =?

OpenStudy (anonymous):

What i do is just pretend to have a sheet of paper and a calculator in your head.

OpenStudy (anonymous):

um..10?

OpenStudy (anonymous):

yeah but im loosing into numbers

OpenStudy (konradzuse):

yup, carry the 1 +9+3?

OpenStudy (konradzuse):

just get used to looking at numbers adding the back to the front.

OpenStudy (anonymous):

im like 2+2=4...4...what?2...4...2....damn..

OpenStudy (raden_zaikaria):

199 + 5431 ....... TAKE 1 AND ADD TO 199 + 1 = 200 AND ADD WITH 5430 ....... 5630 ...

OpenStudy (anonymous):

well thanks :D

OpenStudy (anonymous):

its easier than i was doin

OpenStudy (raden_zaikaria):

BECAUSE OUR BRAIN JUST EASY TO MEMORIZE... 200, 100, 300,

OpenStudy (konradzuse):

pfft don't give the easy way :p Make em work that brain :).

OpenStudy (raden_zaikaria):

yup hehehe our brain is most powrfull than computer...... avoid using calcultor if u can .... practice make perfect hehehe